Circular No. 6792

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science)

                URL http://cfa-www.harvard.edu/iau/cbat.html

Phone 617-495-7244/7440/7444 (for emergency use only)

GRB 971214 L. A. Antonelli, BeppoSAX, Rome; R. C. Butler, Agenzia Spaziale

Italiana, Rome; L. Piro, Istituto Astrofisica Spaziale, CNR, Rome; and G. Celidonio, A. Coletta, A. Tesseri, and C. De Libero, BeppoSAX, Rome, report on behalf of the BeppoSAX team: “A fast follow-up observation of GRB 971214 was carried out with BeppoSAX 6.67 hr after the initial burst. A previously unknown x-ray source (1SAX J1156.4+6513) was clearly detected within the WFC error circle (IAUC 6787) in the LECS and MECS x-ray telescopes, at α = 11^{h}56^{m}25^{s}, δ = +65^{o}13′11′′ (equinox 2000.0; uncertainty ∼ 1′, currently dominated by systematics in the position reconstruction with the new 1-gyro mode). The position of the optical transient (IAUC 6788) can be considered as consistent with the x-ray source. The average source countrate in the range 1.3–10 keV was 0.0056 ± 0.0005 counts/s in two MECS units and 0.0036 ± 0.0005 counts/s in the LECS (0.1–5 keV). This corresponds to a 2–10-keV flux of 4 × 10^{−13} erg cm^{−2} s^{−1}. From a preliminary analysis of the first day of observation, the source faded by a factor of ∼ 5. These results strongly suggest that 1SAX J1156.4+6513 is the x-ray afterglow of GRB 971214.”

M. R. Garcia, A. Muench, and E. Tollestrup, Harvard-Smithsonian Center for Astrophysics; and P. J. Callanan and J. McCarthy, University College, Cork, report on observations made with the Whipple Observatory 1.2-m telescope (+ STELIRCAM) in bands J and K during Dec. 15.519– 15.562 and 16.524–16.569 UT (limiting mag K = 18.5, J = 20.0), in which the K-band and Dec. 16 J-band images cover the entire revised (1′-radius) error circle reported by Antonelli et al., above: “We find a limit of 1 mag to the variability of the dozen stars seen within the error circle; this upper limit is applicable at the position of the fading optical counterpart reported by Halpern et al. (IAUC 6788). Our images suffer due to light scattered from HR 103690, which covers about half of the revised error box.”

COMET C/1997 T1 (UTSUNOMIYA) Visual m_{1} estimates: Nov. 26.74 UT, 9.9 (B. H. Granslo, Movatn, Oslo,

Norway, 0.07-m refractor); Dec. 2.99, 9.9 (J. Bortle, Stormville, NY, 0.41-m reflector); 6.08, 10.5 (C. E. Spratt, Victoria, BC, 0.20-m reflector); 16.69, 9.8 (M. Lehký, Hradec Králové, Czech Rep., 25×100 binoculars).

1997 December 19  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1997 CBAT  Daniel W. E. Green